1. A monopoly’s total cost function is TC = 200 + 8Q + 4Q2. The inverse demand function is P = 400 – 10Q. What will be the monopoly’s profit if it charges a single price to all customers?   

Group of answer choices

a.$2,150

b.$3,420

c.$3,640

d.$2,544

$1,980

2. A Cournot oligopoly has four firms in the industry. The market price elasticity of demand is –2.5 and the marginal cost of production is $200. What is the profit-maximizing price, rounded to the nearest dollar?

Group of answer choices

a.More information is needed to answer this question.

b.$208

c.$500

d.$222

e.$354

3.Assume that a monopoly’s price elasticity of demand is –2.8. If the firm’s marginal cost is $36, what price should the firm charge in order to maximize profit?

Group of answer choices

$56

$72

$42

$64

$90
Answer #1

(1) TC = 200+86 +4Q² d [T] =d (200 +884494]: 8+89 do do MC = Po 400-100 TRE PO = 400 g - 100? MR=[TR]: d (4000-100] -400-200

HOMEWORKLIB POLICY says only 1st question should be answered when multiple questions are posted. Please upvote the solution